sonia is four years old.her elder sister aliya is three time as old as sonia. How old will aliya be when she is twice as old as sonia?
Answer by CubeyThePenguin(3113)

Right now: Sonia is 4, Aliya is 12

2(4+x)= 12 + x
8 + 2x = 12+x
x = 4

In 4 years, Sonia will be 8 and Aliya will be 16.